OpenStudy (anonymous):

How is 10/(1+e^t) supposed to be integrated from limits 0 to 2? Please show steps.

OpenStudy (anonymous):

put 1 +e^t= z diff. w.r.t. z , we get e^t dt= dz

OpenStudy (anonymous):

what is w. and r.?

OpenStudy (anonymous):

ok lets do this way..

OpenStudy (anonymous):

divide the num and den. by e^t/2

OpenStudy (anonymous):

lol, w.r.t means with respect to.

OpenStudy (anonymous):

Sorry, I know you're trying to use substitution, but probably a different method from what I was taught

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

did it help?

OpenStudy (anonymous):

What's the point of dividing num. and den. by e^t/2 though?

OpenStudy (anonymous):

That was another method of doing it

OpenStudy (anonymous):

I do not think dividing by E^t/2 is a correct step in integrating this though

OpenStudy (anonymous):

do with the first method then

OpenStudy (anonymous):

My problem is actually, I know that 1/(1+e^t) is lnl1+e^tl, but I don't know what to do with the 10

OpenStudy (anonymous):

Using substitution doesn't help

OpenStudy (anonymous):

1/(1+e^t) is lnl1+e^tl ???? how did u get that??

OpenStudy (anonymous):

dhatraditya, do you have any input?

OpenStudy (anonymous):

This is integration....so I need to find anti-derivative

OpenStudy (anonymous):

but ur retriceumption is false......1/(1+e^t) is lnl1+e^tl??? no way

OpenStudy (anonymous):

So what is the correct integration of 1/1+e^t; could you show me with steps?

OpenStudy (anonymous):

the first method works for sure......after putting 1+e^t=z u will get this : intergartion (dz/ (z-1)z

OpenStudy (anonymous):

you have yo change your limits appropriately

OpenStudy (anonymous):

it is indefinite integral...therefore no lower n upper limits

OpenStudy (anonymous):

I actually don't care about the limits, I don't know how to antidifferentiate this

OpenStudy (anonymous):

zuuto: will u please substitue what I said.....this works ....I've just done it on paper

OpenStudy (anonymous):

limits are 0 to 2 according to the problem

OpenStudy (anonymous):

ah...my bad

OpenStudy (anonymous):

ankur is right, use substitution. It is the easiest method for definite integrals.

OpenStudy (anonymous):

わかて、 すみません

OpenStudy (anonymous):

oooh. Is that chinese? Sorry I dont understand it.

OpenStudy (anonymous):

sorry, my intention was to understand how to differentiate this, I'm trying it right now

OpenStudy (anonymous):

It was a mistake to type it, it's japanese

OpenStudy (anonymous):

oh i see.

OpenStudy (anonymous):

dont forget to change the limits like dhatraditya said....u can't ignore them when u have integral in z

OpenStudy (anonymous):

Can I make a suggestion? You might want to recognise\[\frac{1}{1+e^t}=\frac{1+e^t-e^t}{1+e^t}=\frac{1+e^t}{1+e^t}-\frac{e^t}{1+e^t}=1-\frac{e^t}{1+e^t}\]You can integrate that directly,\[t-\log (1+e^t)\]

OpenStudy (anonymous):

Sub. in your limits and multiply by 10.

OpenStudy (anonymous):

lokisan is right. His method is superior. and faster.

OpenStudy (anonymous):

period :D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

coincidentally i just figured that out too

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

making the integral into two parts

OpenStudy (anonymous):

so it becomes t-lnle^t+1l I think

OpenStudy (anonymous):

but u need to get used to substitution as well....u can't just rely on making two integrals every time

OpenStudy (anonymous):

Most cases substitution is required, but I'm not how substitution works in this

OpenStudy (anonymous):

yes, substitution is a general method. Splitting was more elegant in this case and lokisan correctly spotted that.

OpenStudy (anonymous):

I'd say now u have the answer ...try with subs. too

OpenStudy (anonymous):

or yeah in the second part of the intergral where you have to integrate e^t/e^t=1, substitution works definitely

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ahhh thanks so muchh

OpenStudy (anonymous):

now I just need to apply this to the problem, which was 10/1+e^t